# Sparse Hanson-Wright Inequalities with Applications

### v2

@@ Abstract
We derive Hanson-Wright inequalities for the quadratic form of a random vector with sparse independent components. Specifically, we consider cases where the components of the random vector are sparse $\alpha$-subexponential random variables with $\alpha>0$. Our proof relies on a novel combinatorial approach to estimate the moments of the random quadratic form. In addition, we obtain a new Bernstein-type inequality for the sum of independent sparse $\alpha$-subexponential random variables. We present two applications with the sparse Hanson-Wright inequality: (1) Local law and complete eigenvector delocalization for sparse $\alpha$-subexponential Hermitian random matrices; (2) Concentration of the Euclidean norm for the linear transformation of a sparse $\alpha$-subexponential random vector.
@@ 1 Introduction
This paper aims to derive Hanson-Wright inequalities for random variables with a sparse structure. We specifically investigate the case where each $X_{i}=x_{i}\xi_{i}$, with $x_{i}\sim\mathrm{Ber}(p_{i})$ and all $x_{i},\xi_{i}$ being independent. The $\xi_{i}$’s are centered $\alpha$-subexponential random variables, characterized by a parameter $\alpha>0$ such that for any $t>0$, $\mathbb{P}(|\xi_{i}|\geq t)\leq 2\exp(-t^{\alpha}/K_{1}^{\alpha})$. Equivalent formulations for $\alpha$-subexponential random variables exist using the Orlicz norm $\|\xi_{i}\|_{\Psi_{\alpha}}<\infty$ or the $L_{r}$ norms: $\|\xi_{i}\|_{L_{r}}\leq K_{2}r^{1/\alpha}$. Note that $K_{1}$, $K_{2}$, and $\|\xi_{i}\|_{\Psi_{\alpha}}$ can be chosen to differ only by a constant factor (depending on $\alpha$). The class of $\alpha$-subexponential random variables includes sub-gaussian ($\alpha=2$), sub-exponential ($\alpha=1$), and heavier-tailed distributions ($\alpha\in(0,1)$). We allow $\alpha=\infty$, which corresponds to bounded random variables satisfying $|\xi_{i}|\leq K$ almost surely (or equivalently, $\|\xi_{i}\|_{L_{r}}\leq K$ for all $r\geq 1$). Our focus is on the sparse regime where $p_{i}\ll 1$.
@@
Throughout this paper, $C,C^{\prime},c$ denote absolute constants that may vary in different equations, and $C_{\alpha}$ is an absolute constant depending only on $\alpha$.
@@
Notations: Let $A=(a_{ij})$ be an $n\times n$ real matrix. The Frobenius norm of $A$ is defined as $\|A\|_{F}=\sqrt{\sum_{i,j}a_{ij}^{2}}$, and its operator norm is denoted by $\|A\|$. The maximum norm is given by $\|A\|_{\max}=\max_{i,j}|a_{ij}|$. We also denote $\|A\|_{1,\infty}=\max_{i}\sum_{j}|a_{ij}|$ and $\|A\|_{2,\infty}=\max_{i}(\sum_{j}|a_{ij}|^{2})^{1/2}$. For a vector $a\in\mathbb{R}^{n}$, $\|a\|$ denotes its Euclidean norm and $\|a\|_{\infty}$ denotes the maximum absolute value of its entries.
For a real random variable $X$, its $L_{r}$ ($r\geq 1$) norm is denoted by $\|X\|_{L_{r}}=(\mathbb{E}|X|^{r})^{1/r}$. The $\alpha$-Orlicz norm [^1] This is a quasi-norm for $\alpha\in(0,1).$ of a random variable $X$ is defined by
@@
For two functions $f(n),g(n)>0$, we use the asymptotic notations $g(n)\ll f(n)$ and $g(n)=o(f(n))$ if $f(n)/g(n)\to\infty$ as $n\to\infty$. Additionally, we use $f(n)\lesssim g(n)$ to denote that $f(n)\leq Cg(n)$ for some $C>0$.
@@
Let $X=(X_{1},\cdots,X_{n})^{T}$ be a random vector with independent centered sparse sub-exponential entries, i.e., $X_{i}=x_{i}\xi_{i}$ where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable satisfying $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$ with some $K>0$. Let $A$ be an $n\times n$ diagonal-free matrix. Then there exists an absolute constant $C>0$ such that for any $t>0$,
$\mathbb{P}(|X^{T}AX|\geq t)\leq C^{\prime}\exp\left(-C\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{1/2}\right\}\right),$ (1.2)
where
$\gamma_{1,\infty}:=\max_{i}\left\{\sum_{j\neq i}|a_{ij}|p_{j},\sum_{j\neq i}|a_{ji}|p_{j}\right\}.$ (1.3)
REMARK 1.1.
We now discuss some generalizations of our results. The assumption that each $X_{i}$ is a product of independent Bernoulli and (centered) sub-exponential random variables can be relaxed in multiple ways. First, the conclusion of Theorem 1 holds for any independent centered random variables $X_{i}$ whose moments satisfy $\mathbb{E}|X_{i}|^{r}\leq p_{i}(Kr)^{r}$ for $r\geq 2$ and some $K>0$, where $p_{i}$ represents the sparsity parameter. Second, with a minor modification of our proof, the theorem extends to independent centered Bernoulli random variables, i.e., $X_{i}=x_{i}-p_{i}$ where $x_{i}\sim\mathrm{Ber}(p_{i})$. In this case, (1.2) holds with $K=1$. Finally, Theorem 1 generalizes to independent noncentered random variables $Y_{i}$, providing concentration results for $\sum_{i\neq j}a_{ij}Y_{i}Y_{j}$ around its mean $\sum_{i\neq j}a_{ij}\mathbb{E}(Y_{i})\mathbb{E}(Y_{j})$. This extension is detailed in Theorem 13 of Appendix B.
@@
$\mathrm{Var}(X^{T}AX)=\sum_{i\neq j}(a_{ij}^{2}+a_{ij}a_{ji})p_{i}p_{j}\mathbb{E}(\xi_{i}^{2})\mathbb{E}(\xi_{j}^{2})\leq 32K^{4}\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}.$
The inequality above follows from $a_{ij}a_{ji}\leq(a_{ij}^{2}+a_{ji}^{2})/2.$ The sub-exponential tail dominates when $A$ has a prominent row or column. For instance, consider a case where only the first row of $A$ is non-zero, with non-negative entries for simplicity, and $X_{i}=x_{i}-p$ for all $i$ where $x_{i}\sim\mathrm{Ber}(p)$. Then $X^{T}AX=X_{1}\cdot(\sum_{i=2}^{n}a_{1i}X_{i})$ is the product of two independent random variables. If $p$ is chosen such that $p(\sum_{i=2}^{n}a_{1i})=p\|A\|_{1,\infty}=O(1)$, then $\sum_{i=2}^{n}a_{1i}x_{i}$ is approximately Poisson with parameter $p\|A\|_{1,\infty}$. The tail behavior of $X^{T}AX$ is controlled by $\exp(-ct/p\|A\|_{1,\infty})$. The $1/2$-subexponential tail emerges due to individual terms in the summation $X^{T}AX=\sum_{i\neq j}a_{ij}X_{i}X_{j}$. Conditioning on $x_{i}$’s, each $a_{ij}X_{i}X_{j}$ is a product of two independent sub-exponential random variables and has a tail of magnitude $\exp(-(t/K^{2}|a_{ij}|)^{1/2})$.
Moreover, for sparse $\alpha$-subexponential random vectors, we have the following generalization of Theorem 1.
THEOREM 2.
Consider a random vector $X=(X_{1},X_{2},\dots,X_{n})^{T}$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$, for any $r\geq 1$ and some $\alpha>0$. Assume $A$ is a diagonal-free $n\times n$ matrix. Then there exists a constant $C_{\alpha}>0$ depending only on $\alpha$ such that for any $t>0$,
$\mathbb{P}(|X^{T}AX|\geq t)\leq C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}\right).$ (1.4)
REMARK 1.2.
We see the third term in the tail bound of Theorem 2 does not gain any improvement for $\alpha>1$. It is an open question to see if Theorem 2 can be improved for $\xi_{i}$ being sub-gaussian. For sparse sub-gaussian random vectors, compared to (1.4), the inequality in [47] (presented later as Theorem 7) has $t/(K^{2}\|A\|)$ in the second term, which is independent of the sparsity $p_{i}$, but it does not involve the $t^{1/2}$ term.
@@
Consider a random vector $X=(X_{1},X_{2},\dots,X_{n})$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable with $\|\xi_{i}\|_{r}\leq Kr^{\frac{1}{\alpha}}$, for all $r\geq 1$ and some $\alpha>0$. Let $a=(a_{1},\ldots,a_{n})$ be a deterministic vector in $\mathbb{R}^{n}$. Then for any $t>0$,
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{i}X_{i}\right|\geq t\right)\leq C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{2}\sum_{i}a_{i}^{2}p_{i}},\left(\frac{t}{K\|a\|_{\infty}}\right)^{\min\{\alpha,1\}}\right\}\right).$ (1.5)
Theorem 3 can be seen as a generalization of Bernstein’s inequality for $\alpha$-subexponential random variables with a dependence on the variance of the linear form. We present a brief discussion of related literature. Directly applying existing results from [22, Corollary 6.1], one will only get $K^{2}\sum_{i}a_{i}^{2}$ in the first denominator of (1.5) without the dependence on $p_{i}$. For sums of independent $\alpha$-subexponential random variables (termed sub-Weibull of order $\alpha$ in their work), [28, Theorem 3.1] established concentration bounds that, while not reflecting the sparsity $p_{i}$ in the sub-gaussian tail, achieve a sharper second term $\exp(-(t/KL_{\alpha}(a))^{\alpha})$, where $L_{\alpha}(a)=\|a\|_{\infty}$ if $\alpha\leq 1$ and $L_{\alpha}(a)=\|a\|_{\frac{\alpha}{\alpha-1}}$ if $\alpha>1$. Regarding variance-scaled tail bounds, [28, Theorem 3.2-3.3] proved the following tail bound:
$2\exp\left(-\min\left\{C\frac{t^{2}}{\sum_{i}a_{i}^{2}p_{i}\mathbb{E}\xi_{i}^{2}},C_{\alpha}\left(\frac{t}{K\|a\|_{\infty}\log^{\frac{1}{\alpha}}(n+1)}\right)^{\min\{\alpha,1\}}\right\}\right).$
The factor $(\log(n+1))^{1/\alpha}$ is shown to be sharp by [1, Section 2.2], assuming only $\|X_{i}\|_{\psi_{\alpha}}<\infty$ and variance-based norm bounds. In contrast, our new bound (1.5) captures the sparsity parameters $p_{i}$ in the sub-gaussian tail with a variance proxy $K^{2}\sum_{i}a_{i}^{2}p_{i}$, while maintaining a dimension-free second term.
We now bound the concentration of the diagonal sum $S_{\mathrm{diag}}$ in the quadratic form $\sum_{i,j}a_{ij}X_{i}X_{j}$ by applying Theorem 3 to $X_{i}^{2}-EX_{i}^{2}$, which are sparse $\frac{\alpha}{2}$-subexponential random variables.
COROLLARY 1.
Consider a random vector $X=(X_{1},X_{2},\dots,X_{n})$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable with $\|\xi_{i}\|_{r}\leq Kr^{\frac{1}{\alpha}}$, for all $r\geq 1$ and some $\alpha>0$. Then for any $t>0$,
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{ii}(X_{i}^{2}-\mathbb{E}X_{i}^{2})\right|\geq t\right)$ (1.6)
$\leq$ $\,C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{4}\sum_{i}a_{ii}^{2}p_{i}},\left(\frac{t}{K^{2}\max_{i}|a_{ii}|}\right)^{\min\{\frac{\alpha}{2},1\}}\right\}\right).$ (1.7)
Combining the bounds for the diagonal and off-diagonal terms, we have the following result for random quadratic forms of a general matrix $A$:
COROLLARY 2.

### v3

@@ Abstract
We derive new Hanson-Wright-type inequalities tailored to the quadratic forms of random vectors with sparse independent components. Specifically, we consider cases where the components of the random vector are sparse $\alpha$-subexponential random variables with $\alpha>0$. When $\alpha=\infty$, these inequalities can be seen as quadratic generalizations of the classical Bernstein and Bennett inequalities for sparse bounded random vectors. To establish this quadratic generalization, we also develop new Bersntein-type and Bennett-type inequalities for linear forms of sparse $\alpha$-subexponential random variables that go beyond the bounded case $(\alpha=\infty)$. Our proof relies on a novel combinatorial method for estimating the moments of both random linear forms and quadratic forms.
We present two key applications of these new sparse Hanson-Wright inequalities: (1) A local law and complete eigenvector delocalization for sparse $\alpha$-subexponential Hermitian random matrices, generalizing the result of He et al. (2019) beyond sparse Bernoulli random matrices. To the best of our knowledge, this is the first local law and complete delocalization result for sparse $\alpha$-subexponeitial random matrices down to the near-optimal sparsity $p\geq\frac{\mathrm{polylog}(n)}{n}$ when $\alpha\in(0,2)$ as well as for unbounded sparse sub-gaussian random matrices down to the optimal sparsity $p\gtrsim\frac{\log n}{n}.$ (2) Concentration of the Euclidean norm for the linear transformation of a sparse $\alpha$-subexponential random vector, improving on the results of Götze et al. (2021) for sparse sub-exponential random vectors.
@@ 1 Introduction
The Hanson-Wright inequality has become a fundamental tool in the study of high-dimensional probability [59], random matrix theory [60], and signal processing [39]. However, The inequality (1.1) is limited to sub-gaussian vectors and it does not capture the variance dependence of $X$. If we consider a random vector with centered Ber$(p)$ entries where $p\to 0$ as $n\to\infty$, the inequality (1.1) holds with $K\asymp|\log p|^{-1/2}$ [50]. Given a Bernoulli random vector $X$, the dependence on $p$ of a linear form $\sum_{i=1}^{n}a_{i}X_{i}$ is well-captured by Bennett’s inequality (or Chernoff bound), which plays a fundamental role in various applications in random graph theory and theoretical computer science [4].
These observations naturally lead to the following question: Is there a generalization of Bennett’s inequality for the quadratic forms of sparse random vectors?
In this paper, we aim to derive Hanson-Wright inequalities for random variables with a sparse structure, and our new concentration inequalities capture the additional log factors analogous to the improvement from Bernstein’s inequality to Bennett’s inequality for linear forms of random vectors.
A canonical example takes the form $X_{i}=x_{i}\xi_{i}$, where $x_{i}\sim\mathrm{Ber}(p_{i})$ and all $x_{i},\xi_{i}$ being independent. The $\xi_{i}$’s are $\alpha$-subexponential random variables (also known as sub-Weibull of order $\alpha$), characterized by a parameter $\alpha>0$ such that for any $t>0$, $\mathbb{P}(|\xi_{i}|\geq t)\leq 2\exp(-t^{\alpha}/K_{1}^{\alpha})$. These random variables can be equivalently defined using either the Orlicz norm $\|\xi_{i}\|_{\Psi\alpha}<\infty$ or the $L_{r}$ norms: $\|\xi_{i}\|_{L_{r}}\leq K_{2}r^{1/\alpha}$. The constants $K_{1}$, $K_{2}$, and $\|\xi_{i}\|_{\Psi_{\alpha}}$ differ only by factors depending on $\alpha$. It can be readily verified that the moments of $X_{i}$ satisfy $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(K_{2}r^{\frac{1}{\alpha}})^{r}$.
The class of $\alpha$-subexponential random variables includes sub-gaussian ($\alpha=2$), subexponential ($\alpha=1$), and heavier-tailed distributions ($\alpha\in(0,1)$). We also allow $\alpha=\infty$, which corresponds to bounded random variables satisfying $|\xi_{i}|\leq K$ almost surely (or equivalently, $\|\xi_{i}\|_{L_{r}}\leq K$ for all $r\geq 1$). The concentration behavior of $\alpha$-subexponential random variables exhibits a transition at $\alpha=1$, with $\alpha\geq 1$ corresponding to log-concave tails and $\alpha\leq 1$ to log-convex tails. Our primary interest lies in the sparse regime where $p_{i}\ll 1$.
More generally, we study sparse $\alpha$-subexponential random variables, which we characterize through their moments:
DEFINITION 1.1 (Sparse $\alpha$-subexponential random variables).
A random variable $X$ is said to be a sparse $\alpha$-subexponential random variable for some $\alpha\in(0,\infty]$ if
$\mathbb{E}(|X|^{r})\leq p\left(Kr^{\frac{1}{\alpha}}\right)^{r}$ (1.2)
for all $r\geq 1$ and some $K>0$, where $p\in(0,1]$ is the sparsity parameter.
Random vectors with sparse $\alpha$-subexponential entries arise in various areas, including large deviation theory [29], random matrix theory [7], and random weighted networks [64], where their concentration properties are of critical importance.
@@
Throughout this paper, $C,C^{\prime},c$ denote absolute constants that may vary in different equations, and $C_{\alpha},c_{\alpha},c_{\alpha}^{\prime}$ are absolute constants depending only on $\alpha$.
@@
Notations: Let $A=(a_{ij})$ be an $n\times n$ real matrix. The Frobenius norm of $A$ is defined as $\|A\|_{F}=\sqrt{\sum_{i,j}a_{ij}^{2}}$, and its operator norm is denoted by $\|A\|$. The maximum norm is given by $\|A\|_{\max}=\max_{i,j}|a_{ij}|$. We also denote $\|A\|_{1,\infty}=\max_{i}\sum_{j}|a_{ij}|$ and $\|A\|_{2,\infty}=\max_{i}(\sum_{j}|a_{ij}|^{2})^{1/2}$. For a vector $a\in\mathbb{R}^{n}$, $\|a\|$ denotes its Euclidean norm and $\|a\|_{\infty}$ denotes the maximum absolute value of its entries. For a real random variable $X$, its $L_{r}$ ($r\geq 1$) norm is denoted by $\|X\|_{L_{r}}=(\mathbb{E}|X|^{r})^{1/r}$. The $\alpha$-Orlicz norm [^1] This is a quasi-norm for $\alpha\in(0,1).$ of a random variable $X$ is defined by
@@
For two functions $f(n),g(n)>0$, we use the asymptotic notations $g(n)\ll f(n)$ and $g(n)=o(f(n))$ if $f(n)/g(n)\to\infty$ as $n\to\infty$. Additionally, we use $f(n)\lesssim g(n)$ or $g(n)\gtrsim f(n)$ to denote that $f(n)\leq Cg(n)$ for some $C>0$. Let $[\log(x)]_{+}$ denote the function equal to $\log(x)$ if $x>1$ and 00 otherwise. In the latter case, we interpret $\frac{1}{[\log(x)]_{+}}=\infty$. For a set $S$, let $\mathbf{1}(S)$ denote the indicator function of $S$.
Below we present our new results on the concentration of the quadratic form for sparse random vectors. We first state Bernstein-type inequalities for random quadratic forms and linear forms in Section 1.1, which is easier to state and serves as an important immediate result towards a Bennett-type inequality for the quadratic form, which we will introduce in Section 1.2.
1.1 Bernstein-type inequalities for sparse random vectors
@@
Let $X=(X_{1},X_{2},\ldots,X_{n})^{T}$ be a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(0,\infty]$. Le $A$ be a diagonal-free $n\times n$ matrix. Then for any $t>0$,
$\mathbb{P}(|X^{T}AX|\geq t)\leq e^{2}\exp\left(-c_{\alpha}\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}\right),$ (1.3)
where $c_{\alpha}$ is an absolute constant depending on $\alpha$ and
$\gamma_{1,\infty}:=\max_{i}\Big{\{}\sum_{j\neq i}|a_{ij}|p_{j},\sum_{j\neq i}|a_{ji}|p_{j}\Big{\}}.$ (1.4)
@@
$\mathrm{Var}(X^{T}AX)=\sum_{i\neq j}(a_{ij}^{2}+a_{ij}a_{ji})\mathbb{E}(X_{i}^{2})\mathbb{E}(X_{j}^{2})\leq 2^{1+\frac{4}{\alpha}}K^{4}\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}.$
The inequality above follows from $a_{ij}a_{ji}\leq(a_{ij}^{2}+a_{ji}^{2})/2.$ The subexponential tail dominates when $A$ has a prominent row or column. For instance, consider a case where only the first row of $A$ is non-zero, with non-negative entries for simplicity, and $X_{i}=x_{i}-p$ for all $i$ where $x_{i}\sim\mathrm{Ber}(p)$. Then $X^{T}AX=X_{1}\cdot(\sum_{i=2}^{n}a_{1i}X_{i})$ is the product of two independent random variables. If $p$ is chosen such that $p(\sum_{i=2}^{n}a_{1i})=p\|A\|_{1,\infty}=O(1)$, then $\sum_{i=2}^{n}a_{1i}x_{i}$ is approximately Poisson with parameter $p\|A\|_{1,\infty}$. The tail behavior of $X^{T}AX$ is controlled by $\exp(-ct/p\|A\|_{1,\infty})$. For $\alpha\in(0,1]$, the $\alpha/2$-subexponential tail emerges due to individual terms in the summation $X^{T}AX=\sum_{i\neq j}a_{ij}X_{i}X_{j}$. Conditioning on $x_{i}$’s, each $a_{ij}X_{i}X_{j}$ is a product of two independent $\alpha$-subexponential random variables and has a tail of magnitude $\exp(-(t/K^{2}|a_{ij}|)^{\alpha/2})$. For $\alpha>1$, while the $1/2$-subexponential tail holds (as these random variables are subexponential), this tail bound is sub-optimal. Subsequently, we derive a sharper tail bound that improves this result (given in Theorem 3).
@@
Let $X=(X_{1},X_{2},\ldots,X_{n})$ be a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(0,\infty]$. Let $a=(a_{1},\ldots,a_{n})$ be a deterministic vector in $\mathbb{R}^{n}$. Then for any $t>0$,
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{i}X_{i}\right|\geq t\right)\leq e^{2}\exp\left(-c_{\alpha}^{\prime}\min\left\{\frac{t^{2}}{K^{2}\sum_{i}a_{i}^{2}p_{i}},\left(\frac{t}{K\|a\|_{\infty}}\right)^{\min\{\alpha,1\}}\right\}\right).$ (1.5)
Theorem 2 can be seen as a generalization of Bernstein’s inequality for $\alpha$-subexponential random variables with a dependence on the variance of the linear form. Our proof technique relies on estimating all moments of the linear form $\langle a,X\rangle$. A discussion of related results to Theorem 2 is given in Section 2.
1.2 Bennett-type inequalities for sparse random vectors
It is worth noting that for $\alpha>1$, the tail bounds in both Theorem 1 and Theorem 2 are sub-optimal as they do not depend on $\alpha$. To address this limitation, we develop new tail bounds that improve upon both theorems in the large deviation regime for all $\alpha>1$. Below, we present a simplified version of these improved bounds, which has better tail probability estimates compared to Theorems 1 and 2 in the large deviation regime.
THEOREM 3 (Simplified tail bounds: $\alpha>1$).